OpenStudy (anonymous):

Factor this expression completely. a^3 - 1 How do you figure this out??

OpenStudy (zzr0ck3r):

a^3-1 = a^3 - 1^3 a^3-b^3 = (a-b)(a^2+ab+b^2). a=a, b = 1

OpenStudy (zzr0ck3r):

(a-1)(a^2+a+1)

OpenStudy (anonymous):

Thank you very much!!